The magnitude of the electric force between two point charge is initially 180 N. What will the force be if the distance between the two charges is doubled? (A) 45 N, (B) 90 N, (C) 180 N, (D) 360 N, (E) 720 N
Answer #1

Force, F = k q1 q2/r^2 = 180

F' = k 2*q1 2*q2/r^2 = 4 k q1 q2/r^2 = 4F

F' = 180 x 4 = 720 N

Comment in case any doubt please rate my answer...